Elliott's Cross Country Transportation Services has a capital structure with 25% debt at an 8% interest rate. Its beta is 1.4, the risk-free rate is 4%, and the market risk premium is 6%. Elliott's combined federal-plus-state tax rate is 25%. What is Elliott's cost of equity? Do not round intermediate calculations. Round your answer to two decimal places. What is its weighted average cost of capital? Do not round intermediate calculations. Round your answer to two decimal places. What is its unlevered cost of equity? Do not round intermediate calculations. Round your answer to two decimal places.

I found cost of equity is 12.40% and WACC is 10.80%. I am having difficulty finding the unlevered cost of equity.
Answer #1

Cost of Equity = 12.40%

WACC = 10.80%

Debt Equity = 25% / 75%

= 0.33333333333

\small Unlevered Beta = \frac{Levered Beta}{(1+[(1-Tax rate)*\frac{Debt}{Equity}]}

\small Unlevered Beta = \frac{1.4}{(1+[(1-0.25) * 0.33333333333}

= 1.4 / 1.25

= 1.12

Unlevered Beta = 1.12

Unlevered Cost of Equity = R_{f} + beta (R_{m}-R_{f})

= 4% + 1.12 (6%)

= 10.72%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.

We were unable to transcribe this image

We were unable to transcribe this image